23.2 If f(t) = 2 sin t + 3 cost, express f(t) as a single sinusoid and hence determine its

amplitude and phase.



1
Expert's answer
2021-07-29T13:34:23-0400

Solution:

"f(t) = 2 \sin t + 3 \cos t"

Let "f(t)=A \sin(t+\theta)" , where "A" is amplitude and "\theta" is phase.

So, "2 \sin t + 3 \cos t=A \sin(t+\theta)"

"\Rightarrow2 \sin t + 3 \cos t=A [\sin t \cos\theta+\cos t \sin \theta]
\\\Rightarrow2 \sin t + 3 \cos t=A \sin t \cos\theta+A\cos t \sin \theta"

On comparing both sides,

"A\cos\theta=2,A\sin\theta=3
\\\Rightarrow \tan\theta=\dfrac32
\\\Rightarrow \theta=56.31\degree"

Then, "A\cos\theta=2"

"\Rightarrow A\cos 56.31\degree=2
\\ \Rightarrow A=3.605"

Thus, amplitude is 3.605 and phase is "56.31\degree" .